# Generating Fine-Grained Human Motions Using ChatGPT-Refined Descriptions

### v1

@@ 3.2.2 Global Token and Part Tokens
For the text condition, we encode the entire fine-grained description with CLIP [29] as a global token $GL$ of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens $PT_{1:k}$, where k denotes the number of body parts. The global token provides overall information about the motion, while part tokens provide local information about different body parts. This design allows the global token and part tokens to jointly guide the diffusion model in generating human motions.
@@ 4 Experiments
Figure 3: Qualitative results with fine-grained motions. We compare our FG-MDM with MDM [35] and MLD [5]. All three models are trained on HumanML3D. For better visualization, some pose frames are shifted to prevent overlap. Please refer to supplementary materials for more video demos.
Figure 4: Qualitative results with stylized motions. All three models are trained on HumanML3D. Please refer to supplementary materials for more video demos.
@@
Top 1 Top 2 Top 3
Real $0.520^{\pm.003}$ $0.703^{\pm.006}$ $0.799^{\pm.005}$ $0.004^{\pm.000}$ $2.959^{\pm.007}$ $9.412^{\pm.046}$
T2M-GPT [41] $0.269^{\pm.008}$ $0.410^{\pm.009}$ $0.507^{\pm.012}$ $3.910^{\pm.069}$ $5.014^{\pm.060}$ $8.128^{\pm.171}$ $2.578^{\pm.122}$
MLD [5] $0.286^{\pm.002}$ $0.443^{\pm.003}$ $0.552^{\pm.003}$ $2.244^{\pm.053}$ $4.725^{\pm.015}$ $8.773^{\pm.058}$ $\textbf{3.317}^{\pm.078}$
MotionDiffuse [42] $0.308^{\pm.007}$ $0.466^{\pm.008}$ $0.571^{\pm.007}$ $2.755^{\pm.087}$ $\textbf{4.483}^{\pm.031}$ $8.333^{\pm.163}$ $1.667^{\pm.127}$
MDM [35] $0.187^{\pm.009}$ $0.309^{\pm.007}$ $0.399^{\pm.008}$ $4.363^{\pm.265}$ $7.331^{\pm.065}$ $8.432^{\pm.117}$ $3.059^{\pm.027}$
FG-MDM $\textbf{0.308}^{\pm.015}$ $\textbf{0.484}^{\pm.005}$ $\textbf{0.593}^{\pm.003}$ $\textbf{0.663}^{\pm.051}$ $5.649^{\pm.035}$ $\textbf{9.476}^{\pm.077}$ $2.632^{\pm.085}$
Table 1: Quantitative results on HumanML3D. To examine the ability in handling fine-grained descriptions, for all the four control methods, we employ their pre-trained models to generate motions by using our fine-grained description as input.
@@
Top 1 Top 2 Top 3
Real $0.418^{\pm.011}$ $0.633^{\pm.011}$ $0.775^{\pm.009}$ $0.038^{\pm.009}$ $2.749^{\pm.034}$ $10.852^{\pm.194}$
MotionDiffuse [42] $\textbf{0.189}^{\pm.008}$ $\textbf{0.296}^{\pm.016}$ $\textbf{0.376}^{\pm.009}$ $16.136^{\pm.648}$ $\textbf{6.513}^{\pm.104}$ $8.948^{\pm.339}$ $1.024^{\pm.027}$
MDM [35] $0.099^{\pm.013}$ $0.167^{\pm.013}$ $0.235^{\pm.019}$ $16.372^{\pm.855}$ $10.501^{\pm.058}$ $8.758^{\pm.233}$ $\textbf{2.242}^{\pm.998}$
FG-MDM $0.151^{\pm.010}$ $0.269^{\pm.013}$ $0.363^{\pm.008}$ $\textbf{0.344}^{\pm.016}$ $9.351^{\pm.073}$ $\textbf{10.707}^{\pm.082}$ $0.871^{\pm.442}$
Table 2: Quantitative results on KIT. Since T2M-GPT [41] and MLD [5] do not release their pre-trained models on KIT. We thus compare our model with MotionDiffuse and MDM only.
In this section, we first elaborate the datasets, evaluation metrics, and implementation details. We then conduct quantitative experiments to compare FG-MDM with current state-of-the-art approaches. In order to completely evaluate FG-MDM, we further present two ablation studies. To show the generalization capability of our model, we further perform qualitative experiments and a user study to examine FG-MDM’s ability in generating fine-grained and stylized motions.
@@ 4.1.1 Datasets
We utilize the HumanML3D [9] dataset and the KIT [26] dataset to train and evaluate our model. HumanML3D is a recently proposed large-scale dataset of motion-text pairs. It consists of 14,616 motion sequences from the AMASS [21] and HumanAct12 [8] datasets, with multiple ways of describing each motion, resulting in a total of 44,970 text annotations. The KIT dataset, on the other hand, is relatively smaller and contains 3,911 motion sequences along with their corresponding 6,353 text descriptions. For both datasets, we use 80% of the data for training, and use the remaining for test.
We preprocess the 44,970 text descriptions from HumanML3D and 6,353 text descriptions from KIT using ChatGPT-3.5. This extends these descriptions into fine-grained ones for our model training.
@@ 4.1.2 Evaluation Metrics
For quantitative experiments, we employ five evaluation metrics to evaluate our model’s ability in fitting the training data: R Precision, FID, Multimodal Dist, Diversity, and Multimodality. R Precision and Multimodal Dist are used to assess the correlation between generated motions and input text. Diversity and Multimodality are utilized to evaluate the diversity of generated motions. FID measures the difference in feature distribution between generated motions and ground truth in latent space, which is utilized to evaluate the quality of generated motions.
@@ 4.1.3 Implementation Details
In our study, the transformer accepts tokens whose feature dimension is 512 as input. We use 4 attention heads and apply a dropout rate of 0.1. The transformer encoder consists of 8 stacked encoder layers to capture complex relationships and hierarchies in the data. For ChatGPT, we adopt the gpt-3.5-turbo API provided by OpenAI. For text encoding, we employ the frozen CLIP-ViT-B/32 model as the encoder. Our batch size is set to 64. Additionally, we set the diffusion step to 1000. On a single NVIDIA GeForce RTX3090 GPU, it takes about 6 days for training our model.
@@ 4.2 Comparison with Prior Work
To evaluate the performance of FG-MDM in handling fine-grained description-based motion generation, we compare our work with four recent motion generation approaches: T2M-GPT [41], MLD [5], MotionDiffuse [42], and MDM [35]. Note that all these four methods achieve state-of-the-art performance when using short text as input. To examine their ability in handling fine-grained descriptions, for all these four methods, we employ their pre-trained models to generate motions by using our fine-grained description as input. In Table 1 and Table 2, we provide experimental results on the HumanML3D and KIT datasets, respectively. Note that on KIT, T2M-GPT and MLD do not release their pre-trained models. We thus compare our model with MotionDiffuse and MDM only on KIT. For all experiments, We run the evaluation 5 times and “±” indicates the 95% confidence interval.
As reported in both tables, we achieve impressive performance when handling the task of fine-grained description-based motion generation. In particular, FG-MDM leads to much shorter FID than prior state-of-the-art. For example, on KIT, both MotionDiffuse and MDM get long FID around 16, while our model gets a figure shorter than 0.35. This indicates FG-MDM can generate motions closer to real motions than other methods, when using fine-grained description as input.
@@ 4.3 Ablation Study
In order to validate the contribution of different modules, next we conduct two ablation studies. The first study is used to examine the contribution of ChatGPT-Generated fine-grained texts, which is performed by replacing the fine-grained description with the original short text. Therefore, only a single token is adopted to guide the diffusion model. The second study is performed by adopting only the global token to check the contribution of part tokens when fine-grained descriptions are used. In Table 3 and Table 4, we give the results on HumanML3D and KIT, respectively. As observed, ablating either fine-grained descriptions or part tokens leads to poorer performance.
4.4 Generalization Capability Study
Note that FG-MDM aims to not only fit the training data but also generate fine-grained and stylized motions beyond the scope of the dataset. Considering that HumanML3D and KIT include very few fine-grained and stylized motions, we customize some texts that are not in the two datasets to show the effectiveness of FG-MDM in generating fine-grained and stylized motions. In Figure 3 and Figure 4, we show motions generated by MDM [35], MLD [5] and our FG-MDM. Note that for all the three methods, we use models trained on HumanML3D to generate motions. In comparison, our method generates motions that are more consistent with the details described in the slightly complex textual descriptions. Even for difficult stylized textual descriptions, our method is capable of generating motions that correspond to the respective styles. This demonstrates the strong generalization capability of FG-MDM. Please refer to supplementary materials for more video demos.
Figure 5: User study results. For each method, a color bar ranging from blue to red is used to represent the percentage of text-to-motion match levels, with blue indicating the least match and red indicating the most match.
@@ 4.5 User Study
To further examine FG-MDM’s generalization capability, we conduct a user study to evaluate the quality of motions generated by our model based on human visual perception. We customize a total of 40 fine-grained and stylized textual descriptions. With these descriptions, we generate motions by using MDM [35], MotionDiffuse [42] and our FG-MDM. We then recruit 10 users for the study. In each question, participants are asked to rate the degree of matching between the generated motion and the textual description on a scale of 0 to 2. The results are given in Figure 5. Overall, users believe that FG-MDM matches texts much better in generating fine-grained and stylized motions than the other two methods. For example, the proportion of MDM and MotionDiffuse obtaining the most matching scores is only slightly higher than half of FG-MDM.
@@ 5 Conclusion
In this study, we use ChatGPT to perform fine-grained paraphrasing on the textual annotations of HumanML3D and KIT. With these fine-grained descriptions, We explore a Fine-Grained Human Motion Diffusion Model. It utilizes fine-grained descriptions of different body parts to guide the training of a diffusion model. This enables it to learn the generation of fine-grained and stylized motions even outside of the distribution of the training dataset. In the future, we would like to improve the quality of fine-grained annotations of human motions.
[…]
$4.363$ $0.399$ $8.432$
$1.050$ $0.427$ $9.509$
0.663 0.593 9.476
Table 3: Ablation study results on HumanML3D. “F” denotes using ChatGPT-generated fine-grained descriptions. “P” represents adopting part tokens.
[…]
$16.372$ $0.235$ $8.758$
$0.549$ $0.360$ 10.829
0.344 0.383 $10.707$
Table 4: Ablation study results on KIT.

### v2

@@ 3.2.2 Global Token and Part Tokens
For the text condition, we encode the entire fine-grained description with CLIP [35] as a global token $GL$ of the transformer. Apart from this global token, we further encode descriptions of different body parts individually with CLIP as part tokens $PT_{1:k}$, where $k$ denotes the number of body parts. The global token serves as an overall condition to guide the diffusion process. Part tokens provide explicit information for fine-grained control of the movements of each body part. Part tokens effectively make up for the ambiguity of the original description text. It greatly enhances our FG-MDM’s ability to understand in-the-wild text, making it outstanding on zero-shot generation tasks.
@@ 4 Experiments
Figure 3: Qualitative results with unseen motions. We compare our FG-MDM with MDM [41] and MLD [6]. All three models are trained on HumanML3D. For better visualization, some pose frames are shifted to prevent overlap. Please refer to supplementary materials for more video demos.
Figure 4: Qualitative results with unseen stylized motions. All three models are trained on HumanML3D. Please refer to supplementary materials for more video demos.
Table 1: Quantitative results on HumanML3D and HuMMan. The model marked with * indicates that both the ChatGPT-Refined text and the manually annotated text provided by HumanML3D are used during training. Bold text is the best result, underlined text is the second-best result. Zero-shot means that the models are evaluated directly on HuMMan after training on HumanML3D.
Methods HumanML3D HuMMan(zero-shot)
@@
Real $0.002^{\pm.000}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$ $0.032^{\pm.002}$ $23.019^{\pm.042}$ $4.709^{\pm.097}$
MAA [4] $0.774^{\pm.007}$ $8.230^{\pm.064}$
T2M-GPT [47] $\textbf{0.116}^{\pm.004}$ $\underline{3.118}^{\pm.011}$ $\textbf{9.761}^{\pm.081}$ $\textbf{9.631}^{\pm.203}$ $27.582^{\pm.073}$ $5.149^{\pm.145}$
MLD [6] $0.473^{\pm.013}$ $3.196^{\pm.010}$ $\underline{9.724}^{\pm.082}$ $14.970^{\pm.472}$ $27.104^{\pm.024}$ $5.493^{\pm.101}$
MotionDiffuse [48] $0.630^{\pm.001}$ $\textbf{3.113}^{\pm.001}$ $9.410^{\pm.049}$ $30.138^{\pm.712}$ $28.747^{\pm.041}$ $5.357^{\pm.045}$
MDM [41] $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $13.375^{\pm.408}$ $27.689^{\pm.055}$ $5.585^{\pm.089}$
FG-MDM $0.663^{\pm.012}$ $5.649^{\pm.024}$ $9.476^{\pm.068}$ $17.180^{\pm.272}$ $\underline{26.867}^{\pm.030}$ $\underline{5.589}^{\pm.124}$
FG-MDM* $0.618^{\pm.009}$ $5.274^{\pm.048}$ $9.563^{\pm.097}$ $\underline{12.460}^{\pm.330}$ $\textbf{26.814}^{\pm.019}$ $\textbf{5.626}^{\pm.100}$
Table 2: Quantitative results on Kungfu. Zero-shot means that the models are evaluated directly on Kungfu after training on HumanML3D.
Methods Kungfu(zero-shot)
@@
Real $0.133^{\pm.010}$ $22.164^{\pm.041}$ $5.351^{\pm.312}$
MAA [4] - - -
T2M-GPT [47] $\textbf{12.652}^{\pm.429}$ $\underline{25.826}^{\pm.041}$ $5.702^{\pm.428}$
MLD [6] $18.524^{\pm.352}$ $27.182^{\pm.020}$ $5.598^{\pm.356}$
MotionDiffuse [48] $26.363^{\pm.337}$ $26.320^{\pm.035}$ $\textbf{6.117}^{\pm.691}$
MDM [41] $16.396^{\pm.466}$ $26.280^{\pm.095}$ $5.468^{\pm.590}$
FG-MDM $19.340^{\pm.797}$ $26.845^{\pm.052}$ $5.142^{\pm.759}$
FG-MDM* $\underline{15.892}^{\pm.567}$ $\textbf{25.325}^{\pm.035}$ $\underline{5.814}^{\pm.479}$
In this section, we first elaborate the datasets, evaluation metrics, and implementation details in Section 4.1. We then conduct quantitative experiments to compare FG-MDM with current state-of-the-art approaches in Section 4.2. To show the generalization capability of our model, we further perform quantitative experiments, qualitative experiments, and a user study to examine FG-MDM’s ability to generate motions beyond the distribution of training datasets. To evaluate our method comprehensively, we design two additional ablation experiments in Section 4.3. The process of designing prompts will also be elaborated in this section.
@@ 4.1.1 Datasets
We utilize the HumanML3D [9] dataset and the KIT [33] dataset to train and evaluate our model. The HuMMan [5] dataset and Kungfu dataset from the Motion-X dataset [22, 25] are employed to assess the models’ zero-shot performance. HumanML3D is a recently proposed large-scale dataset of motion-text pairs. It consists of 14,616 motion sequences from the AMASS [26] and HumanAct12 [10] datasets, with multiple ways of describing each motion, resulting in a total of 44,970 text annotations. The KIT dataset, on the other hand, is relatively smaller and contains 3,911 motion sequences along with their corresponding 6,353 text descriptions. For both datasets, we use 80% of the data for training and the remaining for testing. Motion-X is a large-scale dataset of whole-body motions and whole-body pose annotations, integrating several existing datasets and additional online videos. For zero-shot testing, we utilize 100% of the HuMMan and Kungfu subsets from it. HuMMan is a multi-modal human dataset, containing 744 motion sequences and their corresponding 744 texts descriptions. Kungfu encompasses many human motions related to martial arts, with a total of 1040 motion sequences and their corresponding 1040 texts descriptions.
We preprocess the 44,970 text descriptions from HumanML3D and 6,353 text descriptions from KIT using ChatGPT-3.5. This preprocessing extends these descriptions into fine-grained ones for our model training.
@@ 4.1.2 Evaluation Metrics
We employ three evaluation metrics for quantitative experiments to evaluate our model’s ability to fit the training data: FID, Multimodal Dist, and Diversity. Multimodal Dist assesses the correlation between generated motions and input text. Diversity is utilized to evaluate the diversity of generated motions. FID measures the difference in feature distribution between generated motions and ground truth in latent space, which is used to evaluate the quality of generated motions.
@@ 4.1.3 Implementation Details
In our study, the transformer accepts tokens whose feature dimension is 512 as input. We use four attention heads and apply a dropout rate of 0.1. The transformer encoder consists of 8 stacked encoder layers to capture complex relationships and hierarchies in the data. For ChatGPT, we adopt the gpt-3.5-turbo API provided by OpenAI. For text encoding, we employ the frozen CLIP-ViT-B/32 model as the encoder. Our batch size is set to 64. Additionally, we set the diffusion step to 1000. On a single NVIDIA GeForce RTX3090 GPU, it takes about six days to train our model.
@@ 4.2 Comparison with Prior Work
To evaluate the performance of FG-MDM in handling zero-shot text-conditioned motion generation, we compare our work with five recent motion generation approaches: MAA [4], T2M-GPT [47], MLD [6], MotionDiffuse [48], and MDM [41]. In Table 1 and Table 2, we provide experimental results on the HumanML3D, HuMMan, and Kungfu datasets, respectively. For all experiments, We run the evaluation five times, and “±” indicates the 95% confidence interval. For the six SOTA methods, on HumanML3D, we directly cite their results reported in their original papers. To examine the generalization ability of the methods, we use HumanML3D as the training set and HuMMan and Kungfu as the test sets. To do so, we train TMR [32, 30] using the HuMMan and Kungfu datasets to obtain a pair of text encoder and motion encoder for calculating the MM Dist metric. For SOTA methods, we apply their released pre-trained models on HumanML3D to HuMMan and Kungfu to evaluate their zero-shot generation performance. Since MAA [4] does not release the pre-trained model, we cannot test its zero-shot generation performance.
When evaluated on the test set of HumanML3D, all five methods achieve state-of-the-art performance. For FG-MDM, the ChatGPT-Refined fine-grained textual description doesn’t match the manually annotated textual description better. Therefore, under within-dataset settings, our model does not exceed those SOTA models on HumanML3D. However, on the HuMMan and Kungfu datasets, FG-MDM captures most of the best and second-best results. Note that the size of our training dataset is much smaller than some SOTA methods like [4], but we still demonstrate solid zero-shot capabilities.
Existing quantitative results do not reflect the real generation performance of the model well. Therefore, we provide some qualitative results to let readers intuitively feel the superiority of our method. In Figure 3 and Figure 4, we show motions generated by MDM [41], MLD [6] and our FG-MDM. Note that for all three methods, we use models trained on HumanML3D to generate motions. In comparison, our method generates motions more consistent with the details described in the fine-grained textual descriptions. This shows that our divide-and-conquer method works. Motion generation models require clear and specific conditions to generate the motions needed. Even for difficult stylized descriptions, our method can generate motions that correspond to the style in the textual descriptions. This demonstrates the strong generalization capability of FG-MDM. Please refer to supplementary materials for more video demos.
Table 3: Ablation study results on HumanML3D. “Fine-grained” denotes using ChatGPT-generated fine-grained descriptions. “Part” represents adopting part tokens. Note that the models are trained on HumanML3D.
[…]
$4.363$ $7.298$ $8.432$
$1.050$ $6.778$ 9.509
0.663 5.649 $9.476$
Table 4: Ablation study results on KIT. Note that the models are trained on KIT.
[…]
$16.372$ $10.502$ $8.758$
$0.549$ $9.826$ 10.829
0.344 9.352 $10.707$
@@ 4.3 Ablation Study
To validate our contribution, we conduct two ablation studies. As shown in Table 3 and Table 4, the first row shows our baseline. The first study examines the contribution of ChatGPT-Generated fine-grained texts, which is performed by replacing the original short text with the fine-grained description. The improvement can be said to be huge. We cleverly utilize the powerful reasoning capabilities of LLMs and let them help our generative model better understand the nature of text conditions, bringing a leap to the zero-shot performance. The second study checks the contribution of part tokens when fine-grained descriptions are used. As observed, a reasonable framework also improves the quality of generated motions. However, perhaps more conditions bring more constraints to generations, leading to a decrease in diversity. But this drop is acceptable. So, we finally adopt the design of part tokens.
The process of designing a prompt is also worth introducing. We put a lot of effort into prompt engineering to better utilize LLMs. Initially, we wanted to replace the abstract emotion words in the text with concrete motion descriptions. The prompt is designed as “Replace the emotional words in this sentence with corresponding descriptions of human body movements, reorganize the language, and output new sentences. [sentence]”. However, GPT cannot understand the task requirements well, and the answers are incorrect. For example, the answers describe muscles, facial expressions, etc., and have nothing to do with human skeletal movements. Then we follow Action-GPT [17] to make the prompt as “Describe a person’s body movements who is performing the action [sentence] in details”. But the output is not concise enough. There would be redundant information and irrelevant information. The structure of the output content is also irregular. For example, The granularity is too fine, and some sentences have no action information. So we decided to divide the body parts to make the output cleaner, i.e., “Use a concise and clear paragraph to describe the body movements of the person executing [sentence] using one paragraph. body movements only includes [‘arms’, ‘legs’, ‘torso’, ‘neck’, ‘buttocks’, ‘waist’].” But it still cannot completely solve this problem. Finally, we add in-context learning to improve output quality. By letting LLMs learn from the examples we give, its output content is standardized, making it better meet our needs. The final prompt is shown in Section 3.1.
Figure 5: User study results. For each method, a color bar ranging from blue to red represents the percentage of text-to-motion match levels, with blue indicating the least match and red indicating the most match.
@@ 4.4 User Study
To further examine FG-MDM’s generalization capability, we conduct a user study to evaluate the quality of motions generated by our model based on human visual perception. We customize a total of 40 textual descriptions beyond the distribution of the dataset. With these descriptions, we generate motions by using MDM [41], MotionDiffuse [48], and our FG-MDM. We then recruit ten users for the study. In each question, participants are asked to rate the degree of matching between the generated motion and the textual description on a scale of 0 to 2. The results are given in Figure 5. Apparently, FG-MDM matches texts much better in generating motions beyond the distribution of the dataset than the other two methods. Nearly half of the generated motions get the highest score. In contrast, MDM and MotionDiffuse perform poorly. Most of the generated motions are not satisfactory.
@@ 5 Conclusion
In this study, we used LLMs to perform fine-grained paraphrasing on the textual annotations of HumanML3D and KIT. With these fine-grained descriptions, We explored a Fine-Grained Human Motion Diffusion Model. It utilizes fine-grained descriptions of different body parts to guide the training of a diffusion model. This enables it to learn the essence of motions and thus generate motions beyond the distribution of training datasets. In the future, we would like to improve the quality of fine-grained annotations of human motions. Having high-quality text labels will greatly promote research on human motion generation.